Fill in the boxes: 
▯ - 5 / 8 = 1 / 4
 Fill in the boxes: 
▯ - 5 / 8 = 1 / 4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

▯- 5 / 8 = 1 / 4
▯=1 / 4 + 5 / 8
▯= [(1 × 2 + 5)] / 8
▯=7 / 8

​Thanks